Currents in dc transmission lines can be 100 A or higher. Some people are concerned that the electromagnetic fields from such lines near their homes could pose health dangers.

Part A

For a line that has current 190 A and a height of 8.0 m above the ground, what magnetic field does the line produce at ground level? Express your answer in teslas.

B =   T  

SubmitMy AnswersGive Up

Part B

What magnetic field does the line produce at ground level as a percent of the earth's magnetic field, which is 0.50 G.

Express your answer using two significant figures.

BBE =

9

  %  

SubmitMy AnswersGive Up

Incorrect; Try Again; 5 attempts remaining

Part C

Is this value of magnetic field cause for worry?

Is this value of magnetic field cause for worry?
No. Since this field is does not differ a lot from the earth's magnetic field, it would be expected to have almost the same effect as the earth’s field.
Yes. Since this field is does not differ a lot from the earth's magnetic field, it would be expected to have almost the same effect as the earth’s field.
Yes. Since this field is much greater than the earth's magnetic field, it would be expected to have more effect than the earth’s field.
No. Since this field is much smaller than the earth's magnetic field, it would be expected to have less effect than the earth’s field.

SubmitMy AnswersGive Up
Answer #1

a)

i = current in the line = 190 A

r = distance of the ground from the line = 8 m

Magnitude of magnetic field at the ground is given as

B = (\mu _{o}/4\pi) (2i/r)

B = (10-7) (2 (190)/8)

B = 4.75 x 10-6 T

b)

Be = Earth's magnetic field at the ground = 0.50 G = 0.50 x 10-4 T = 50 x 10-6 T

B x 100/Be = (4.75 x 10-6) x 100/(50 x 10-6) = 9.5%

c)

No. Since this field is much smaller than the earth's magnetic field, it would be expected to have less effect than the earth’s field.